### v2

# Explainable Molecular Property Prediction: Aligning Chemical Concepts with Predictions via Language Models

Abstract
Providing explainable molecular property predictions is critical for many scientific domains, such as drug discovery and material science. Though transformer-based language models have shown great potential in accurate molecular property prediction, they neither provide chemically meaningful explanations nor faithfully reveal the molecular structure-property relationships. In this work, we develop a new framework for explainable molecular property prediction based on language models, dubbed as Lamole, which can provide chemical concepts-aligned explanations  [^1] Lamole is from the name of a historical winery in Italy called Lamole di Lamole.. We first leverage a designated molecular representation — the Group SELFIES — as it can provide chemically meaningful semantics. Because attention mechanisms in Transformers can inherently capture relationships within the input, we further incorporate the attention weights and gradients together to generate explanations for capturing the functional group interactions. We then carefully craft a marginal loss to explicitly optimize the explanations to be able to align with the chemists’ annotations. We bridge the manifold hypothesis with the elaborated marginal loss to prove that the loss can align the explanations with the tangent space of the data manifold, leading to concept-aligned explanations. Experimental results over six mutagenicity datasets and one hepatotoxicity dataset demonstrate Lamole can achieve comparable classification accuracy and boost the explanation accuracy by up to $14.8\%$, being the state-of-the-art in explainable molecular property prediction.
1 Introduction
Molecular property prediction aims to reveal the relationships between molecular structures and their properties, assisting scientists in screening molecules for various applications such as drug discovery and material design [1, 2, 3, 4, 5]. Several learning-based models are devised based on the underlying molecular representations, such as graph-based molecular representations and string-based molecular representations. Among them, string-based molecular representations, e.g., simplified molecular input line entry systems (SMILES [6]), stand out for their simplicity and adaptability [2, 7, 8]. By viewing the string-based molecular representation as a form of "chemical" language, the Transformer-based language models (LMs) like Bert [9] offer higher throughput and accuracy for molecular property prediction [2, 10].
Figure 1: (a) The molecule visualization of prediction/explanation. The interaction between the benzene ring and the nitro group (highlighted in red) induces the mutagenic property of the molecule. (b)-(e) are the explanation results obtained with various methods: (b) OrphicX [11]; (c) GNNExplainer [12], (d) GNN with gradient-based explainability technique (GradCAM [13]); (e) Bert with GradCAM (molecular string SMILES as input); (f) Bert with GradCAM (molecular string Group SELFIES [8] as the input representation); (g) Our method Lamole assigns an importance score to each functional group/fragment to indicate their contribution to the property.
Despite the superior performance of learning-based prediction methods, what key factors induce the model’s predictions remain largely unexplored, impeding further advancements in the scientific domains. Typically, it is crucial to obtain explanations of predictions while achieving accurate predictions. These obtained explanations could be used for scientific hypotheses validation or/and providing actionable insights for refining investigations, such as optimization for molecular structural design [14, 15, 16]. With different types of molecular representations, explainability techniques for graph neural networks (GNNs) or LMs might be adopted to alleviate the general lack of explainability in molecular prediction [17, 18, 19, 20].
Figure 2: The SMILES and Group SELFIES strings of p-nitrobenzoic acid molecule ($\mathrm{C_{7}H_{5}NO_{4}}$): The tokens in the Group SELFIES string highlighted by color are the corresponding functional groups.
However, we argue that existing explainability techniques often struggle to generate plausible explanations that can highlight chemically meaningful substructures and faithfully uncover the structure-property relationships simultaneously. Specifically, 1) from the molecular representation perspective, the commonly used representations do not explicitly encode the chemically meaningful substructures; current explainability methods can only highlight individual atoms and bonds as explanations (see Fig. 1 (b)$\sim$(e)). 2) From the perspective of explainability techniques, current methods suffer from two main limitations. First, they cannot effectively capture the interactions between functional groups within the molecular structure. Second, they could not generate explanations that align with the domain knowledge of chemists. As a result, they fail to produce explanations that faithfully reflect the structure-property relationships. (see Fig. 1 (f)). Therefore, a novel framework is imperative to provide plausible explanations for molecule property predictions.
The recently proposed string-based molecular representation — Group SELFIES [8] — encodes molecules at the functional group/fragment level, showcasing the possibility of obtaining chemically meaningful explanations. As shown in Fig. 2, Group SELFIES converts a p-nitrobenzoic acid molecule to a string, which explicitly encodes functional groups as tokens, including a benzene, a nitro group, and a carboxyl group. Therefore, with the rich chemical semantics of Group SELFIES, this work develops a new explainable molecular property prediction framework that can provide chemical concepts-aligned explanations based on language models, called Lamole (see Fig. 1 (g)). In particular, we address the above issues with the following contributions.
  1. 1.
    Lamole uses the molecular representation — Group SELFIES — as input tokens for providing the chemically meaningful semantics. By deriving the gradients regarding the input functional group/fragment tokens, gradient-based explanations are generated to indicate the contribution of each functional group/fragment to the molecular properties.
  2. 2.
    By disentangling the information flows of Transformer-based LMs, we observe that the gradient-based explanations cannot well capture the functional group interactions. Therefore, we further integrate attention weights into gradients to generate explanations that can perceive the functional group interactions. The intuition is that the attention weights in the information flows can inherently capture the functional group interactions within the input.
  3. 3.
    To make the explanations more faithfully respect the structure-property relationships, we elaborate on one marginal loss to calibrate the explanations by aligning them with the chemists’ annotations. We show that using only a few molecules with ground truth annotations can significantly improve the explanation accuracy by up to 5%.
  4. 4.
    We first bridge the manifold hypothesis with explainable molecular property prediction. We theoretically demonstrate that the elaborated marginal loss aligns explanations with the data manifold, respecting the structure-property relationship.
Experimental results over six mutagenicity datasets and one hepatotoxicity dataset demonstrate that Lamole can achieve comparable classification accuracy and improve the explanation accuracy by up to $14.8\%$. We also quantitatively evaluate explanations based on the first proposed plausibility metric. Compared to alternative baselines, the explanation plausibility of Lamole is improved by up to $9\%$. Extensive experimental studies demonstrate Lamole achieves state-of-the-art performance in explainable molecular property prediction.
2 Related Work
Explainable Molecular Property Prediction. To gain insights into the model predictions, several explainable GNNs are proposed to explain the relationship between the input molecular graph and the prediction [11, 12, 21, 20]. However, as shown in Fig. 1 (b)$\sim$(d), they only provide explanations at individual atoms and bond levels instead of chemically meaningful substructures. On the other hand, with string-based molecular representations, LMs show great potential in molecular property prediction [10, 22, 23, 24]. However, the "black-box" characteristics of LMs hamper trust and efficient use of these potent computational tools in scientific domains.
Explainability Techniques in LMs. Transformers, lying the cornerstone of LMs, have advanced the state-of-the-art in various research fields, including molecular property prediction [10, 22, 24]. One way to explain their output is to use the attention weights over the input tokens. However, recent studies suggest that "attention is not explanation" because attention weights are often inconsistent with the feature importance [25, 26, 27]. Perturbation-based methods perturb the inputs and evaluate the output changes to reveal the relationship between the input and output. However, the generated explanations may change drastically with very small perturbations [28]. Gradient-based methods determine the feature importance by the partial derivatives of the output to each feature [13]. However, several works show that the gradient-based methods may not be reliable, as they disregard the influence of model architectures on the output and fail to incorporate the information of the model architectures into the explanations [29, 28, 30]. To alleviate this issue, by disentangling the model architectures’ information flows, we combine the attention weights and gradients to generate explanations that faithfully reveal the structure-property relationship.
3 Methodology
Problem Setup. Given a molecular graph $g^{(i)}\in\mathcal{G}=\{(g^{(i)},y^{(i)})\}$, explainable molecular property prediction aims to train a model $f$ to map a molecule $g$ to its property $y$, denoted as $f:g\rightarrow y$, while providing an importance score vector $\mathbf{v}^{(i)}=\{\mathbf{v}_{1}^{(i)},\ldots,\mathbf{v}_{j}^{(i)}\}$ to indicate the contribution of $j$-th functional group/fragment to its property $y$. Particularly, this work proposes to convert the molecular graph $g^{(i)}$ into the Group SELFIES string represented as $s^{(i)}=\{t_{1}^{(i)},\ldots,t_{j}^{(i)}\}$, where $t_{j}^{(i)}$ is the $j$-th functional group/fragment’s token, i.e., [$\cdot$] in Fig. 2. In addition, this work calibrates the explanations in a supervised manner. For this purpose, a few molecules with annotation masks are provided. The annotation masks $\mathbf{m}(g^{(i)})\in\{0,1\}^{j}$ indicate whether $t_{j}^{(i)}$ is the token of a ground truth substructure, where $\mathbf{m}_{j}(g^{(i)})=1$ denotes the substructure corresponding to the token $t_{j}^{(i)}$ inducing the molecular property of $g^{(i)}$. This work uses $\mathcal{D}=\{(s^{(i)},y^{(i)})\}$ with a few annotation masks to learn a model $f$ for explainable molecular property prediction. We omit the superscript (i) for simplicity in the following parts.
Figure 3: An illustration of Lamole. Left panel: Group SELFIES strings are tokenized for fine-tuning the pre-trained language model, and an MLP classifier is equipped with a cross-entropy loss $\mathcal{L}_{CE}$ for molecular property prediction. Right panel: We disentangle the information flows of the Transformer to assert that both attention weights and gradient determine the output. Therefore, we incorporate the attention weights and gradients together to generate importance scores $\mathbf{v}$ as explanations. In addition, a marginal loss $\mathcal{L}_{M}$ is designed to align explanations with the chemists’ annotations $\mathbf{m}$.
3.1 Our Design: Lamole
In this work, we use pre-trained Transformer-based LMs, e.g., Bert family models, as our backbone. Then, we fine-tune the LMs on the Group SELFIES corpus for explainable molecular property prediction. An illustration of proposed Lamole is in Fig. 3. In what follows, we will introduce the detailed design of Lamole.
Fine-Tuning Stage. Fig. 3 shows the fine-tuning stage of the proposed Lamole. We assume that the Transformer encoder in Lamole stacks $L$ identical Transformer layers to encode the molecular string $s$ as token embedding $\mathbf{h}^{(l)}=\{\mathbf{h}_{1}^{(l)},\ldots,\mathbf{h}_{j}^{(l)}\}$, where $\mathbf{h}^{(l)}$ is the token embedding at the $l$-th layer. We use the self-attention weighted average embedding $\mathbf{h}_{o}=(\sum_{j=1}\alpha_{j}\cdot\mathbf{h}_{j}^{(L)})/(\sum_{j=1}\alpha_{j})$ for molecular property prediction, where $\alpha_{j}$ is the attention weight of the $j$-th token. A multilayer perceptron (MLP) classifier is added to predict the molecular property $\hat{y}=\mathrm{MLP}(\mathbf{h}_{o})$ by minimizing the classification cross-entropy loss $\mathcal{L}_{CE}(y,\hat{y})$.
Information Flow-Based Explanations. To derive the explanations from Transformer-based LMs, a common practice is to use gradient-based methods to determine each feature’s importance by analyzing the output’s partial derivatives to each input feature [13]. However, we argue that using gradients alone cannot effectively capture functional group interactions. As depicted in Fig. 1 (f), the gradient-based method CradCAM incorrectly attributes the property to the nitro group and a benzene ring that is not connected to the nitro group.
We illustrate the possible reason by disentangling the information flows of the Transformer. As shown in Fig. 3 right panel, due to the skip connection and attention mechanism, both attention weights, gradients, and the input contribute to the outputs. Therefore, using gradients alone as explanations could fall short of capturing interactions. To address the issue, we leverage both attention and gradients, as well as the input, to derive explanations. Below, we will elaborate on integrating attention weights into gradient-based explanations.
Firstly, we show the process of deriving the gradient-based explanations. Similar to GradCAM [13], the gradient with respect to the $j$-th token’s embedding $\mathbf{h}_{j}^{(l)}$ at the $l$-th layer is derived by
$\nabla\mathbf{h}_{j}^{(l)}=\frac{\partial\hat{y}}{\partial\mathbf{h}_{j}^{(l)}},$ (1)
where $\nabla\mathbf{h}_{j}$ signifies the importance of the $j$-th token in relation to the predicted property $\hat{y}$. Due to the skip connection in Fig. 3 right panel, the input and its corresponding gradient should be leveraged together, and the weighted importance $\mathbf{w}$ of the $j$-th token at the $l$-th layer can be determined by
$\mathbf{w}_{j}^{(l)}=\nabla\mathbf{h}_{j}^{(l)}\odot\mathbf{h}_{j}^{(l)},$ (2)
where $\odot$ is the Hadamard product. The weighted importance is regarded as the gradient-based explanation.
The interaction among tokens can be revealed by the self-attention mechanism in Fig. 3 right panel. The attention mechanism calculates pairwise similarity scores between all pairs of tokens to determine attention weights, and these attention weights inherently encode the functional group interactions. Therefore, we combine the attention weights with the gradient-based explanation to capture functional group interactions. Assuming the attention weights of the $j$-th token at the $l$-th layer is $\alpha_{j}^{(l)}$, we integrate the attention weights with weighted importance $\mathbf{w}$ to consider the interactions. The importance score of the molecule $g$ can be obtained by
$\mathbf{v}_{j}^{(l)}(g)=\left(\tanh(\overline{\alpha_{j}^{(l)}})\cdot\tanh(\mathbf{w}_{j}^{(l)})\right)^{\frac{1}{2}},$ (3)
where $\overline{\alpha_{j}^{(l)}}$ is the averaging of attention weights of multiple attention heads. Finally, we sum $\mathbf{v}_{i}^{(l)}$ over all layers as the final importance score of the $j$-th token,
$\mathbf{v}(g)=\mathrm{softmax}(\sum_{l=1}^{L}\mathbf{v}^{(l)}(g)).$ (4)
The higher the importance scores, the greater the contribution of the corresponding functional groups/fragments to the molecular property.
Towards Plausible Explanations. One plausible explanation should faithfully uncover the structure-property relationships. In other words, the explanation should match the ground-truth substructures with high confidence. Nevertheless, the importance scores of ground-truth substructures might not be significantly higher than those of other parts. To address this issue, we propose a marginal loss to explicitly align explanations with the chemists’ annotations to improve the explanations’ plausibility.
First, we formally define the plausibility of explanations. "Plausibility" refers to how the interpretation convinces humans [31, 32, 33]. Similarly, in our context, "plausibility" refers to the degree of confidence in the explanations that would convince the chemists.
Definition 1
(Plausibility): Given the importance scores $\mathbf{v}$ over all tokens in the molecule $g$, the mean importance score $\mathbf{v}_{\in\mathcal{T}_{g}}$ over ground truth substructures $\mathcal{T}_{g}$ and the mean importance score $\mathbf{v}_{\notin\mathcal{T}_{g}}$ over other substructures $\mathcal{T}_{g}$ are denoted by $\mathbf{v}_{\in\mathcal{T}_{g}}=\frac{\sum_{j}\mathbf{v}_{j}\cdot\mathbb{I}(t_{j}\in\mathcal{T}_{g})}{\sum_{j}\mathbb{I}(t_{j}\in\mathcal{T}_{g})}$ and $\mathbf{v}_{\notin\mathcal{T}_{g}}=\frac{\sum_{j}\mathbf{v}_{j}\cdot\mathbb{I}(t_{j}\notin\mathcal{T}_{g})}{\sum_{j}\mathbb{I}(t_{j}\notin\mathcal{T}_{g})}$, respectively, where $\mathbb{I}(\cdot)$ is the indicator function. The explanations’ plausibility $\mathrm{EP}(g)$ is defined as the ratio of the difference between $\mathbf{v}_{\in\mathcal{T}_{g}}$ and $\mathbf{v}_{\notin\mathcal{T}_{g}}$ to $\mathbf{v}_{\notin\mathcal{T}_{g}}$, $\mathrm{EP}(g)=\frac{\mathbf{v}_{\in\mathcal{T}_{g}}-\mathbf{v}_{\notin\mathcal{T}_{g}}}{\mathbf{v}_{\notin\mathcal{T}_{g}}}.$ (5)
The higher the $\mathrm{EP}$ value, the greater the confidence of the explanation in matching the ground truth substructure.
In our work, the ground truth substructures are annotated by a binary mask vector $\mathbf{m}(g)\in\{0,1\}^{j}$. It is worth noting that using only a few annotations can significantly improve the explanation accuracy. Specifically, the mask vector $\mathbf{m}$ enforces the explanations to align with the ground truth substructures. To achieve the goal, a max-margin loss is designed by maximizing the mean value of the importance scores of tokens that have mask values of 1 while minimizing the mean value of importance scores for tokens with mask values of 0.
$\mathcal{L}_{M}(\mathbf{v},\mathbf{m})=\mathbb{E}_{g\in\mathcal{G}}\left[\max\left(0,\frac{\sum_{j=1}(1-\mathbf{m}_{j}(g))\cdot\mathbf{v}_{j}(g)}{N_{s}}-\frac{\sum_{j=1}\mathbf{m}_{j}(g)\cdot\mathbf{v}_{j}(g)}{N_{c}}\right)+\triangle_{1}\right],$ (6)
where $\triangle_{1}$ is a margin term, $N_{s}$ is the number of tokens with mask values $\mathbf{m}(g)$ of 0, and $N_{c}$ is the number of tokens with mask values $\mathbf{m}(g)$ of 1. The overall optimization objective of the fine-tuning stage is to minimize $\mathcal{L}_{CE}+\mathcal{L}_{M}$. By minimizing $\mathcal{L}_{M}$, Lamole aligns the explanations with the domain knowledge, enhancing the plausibility of explanations.
The next section will theoretically show that by using the designed marginal loss, the explanations can faithfully reflect the structure-property relationships.
3.2 Theoretical Analysis
We bridge the manifold hypothesis with the marginal loss to theoretically show that the explanations can respect the structure-property relationships. Before giving the proof, the notation and definition regarding the manifold hypothesis are presented.
Manifold Hypothesis. It is widely believed that natural data, including molecules, distribute around a manifold [34, 11, 35, 36]. According to the manifold hypothesis for gradient-based explanations [34], if a feature lies in the tangent space of a manifold, then the feature respects the manifold and contributes to the class, and such a feature is desirable to be explained. We call these features "causal features" in our work. Conversely, if a feature is orthogonal to the manifold, then the feature does not contribute to the class. We call these features "spurious features".
With the annotation masks, the causal features $s^{*}$ and spurious features $\overline{s}^{*}$ can be distinguished by $s^{*}=s\odot\mathbf{m}(g)$ and $\overline{s}^{*}=s\odot(1-\mathbf{m}(g))$, respectively, where $\overline{s}^{*}\cup s^{*}=s$ and $s^{*}\cap\overline{s}=\varnothing$. By projecting the causal features and spurious features into the data manifold $\mathcal{M}$, the corresponding manifold regarding the causal features and spurious features can be defined as follows,
Definition 2
(Causal feature manifold and spurious feature manifold): Assume the distribution $p(g|y)$ is implicitly modeled by a manifold $\mathcal{M}$, and the manifold can be decomposed into two components, $\underbrace{p(g|y)}_{\mathcal{M}}=\underbrace{p(g|y)\odot\mathbf{m}(g)}_{\mathcal{M}_{c}}+\underbrace{p(g|y)\odot(1-\mathbf{m}(g))}_{\mathcal{M}_{s}},$ (7) where $\mathcal{M}_{c}$ is the causal feature manifold and $\mathcal{M}_{s}$ represents the spurious feature manifold.
With this decomposition, we demonstrate how the gradient-based explanations $\nabla_{g}\log p(y|g)$ can uncover the structure-property relationships. Due to the page limitation, we provide the proof in Appendix 1.
Theorem 1
The marginal loss of Eq. (6) aligns the gradient-based explanations $\nabla_{g}\log p(y|g)$ with the tangent space of the causal feature manifold $\mathcal{M}_{c}$, thus respecting the structure-property relationships.
4 Experiments
4.1 Experimental Setup
Datasets. We use six datasets on two types of tasks, i.e., hepatotoxicity and mutagenicity, to evaluate the algorithmic performance for the explainable molecular property prediction. Six mutagenicity datasets are Mutag [37], Mutagen [38], PTC-FM [39], PTC-FR [39], PTC-MM [39], and PTC-MR [39]. For hepatotoxicity [39], the Liver dataset [40] is used. We used these datasets as the ground truth substructures in these datasets are known. For detailed information on these datasets, please refer to Appendix 2.
Evaluation Metrics. We use three metrics for evaluating the performance of the proposed Lamole.
  1. 1.
    Classification Accuracy: We evaluate the model’s predictions by $\sum_{i=1}^{I}\mathbb{I}(y^{(i)}=\hat{y}^{(i)})/I$.
  2. 2.
    Explanation Accuracy: We follow the experimental settings in GNNExplainer [12], which formulates the explanation problem as a binary classification of edges. We treat edges inside ground-truth substructure as positive edges and negative otherwise, and AUC is adopted as the metric for quantitative evaluation. We only consider the mutagenic/hepatotoxic molecules because no explicit substructures exist in nonmutagenic/nonhepatotoxic ones.
  3. 3.
    Explanations’ Plausibility: We use the defined explanations’ plausibility $\mathrm{EP}$ to measure how confident the explanation aligns with the ground truth.
Table 1: Mean Classification Accuracy on the Seven Datasets (%)
Methods Mutag Mutagen PTC-FM PTC-FR PTC-MM PTC-MR Liver
GCN [41] 84.6 78.9 54.8 63.0 57.8 53.3 41.1
DGCNN [42] 85.8 74.8 57.3 63.5 61.0 58.6 44.6
edGNN [43] 86.9 75.2 59.8 65.7 64.4 56.3 44.5
GIN [44] 87.5 82.3 62.1 66.2 65.1 64.0 44.9
RW-GNN [45] 87.2 80.3 61.9 64.0 62.4 57.0 43.2
DropGNN [46] 89.4 80.7 62.0 66.0 63.7 64.2 45.0
IEGN [47] 84.6 80.1 60.8 59.8 61.1 59.5 45.3
ChemBERTa [10] 86.8 78.0 60.0 65.7 60.4 58.7 45.7
Lamole+DistilBert 84.2 76.8 57.5 69.0 60.2 64.5 47.2
Lamole+DeBerta 86.8 73.7 58.6 69.5 59.7 63.8 45.8
Lamole+Bert 88.2 74.5 62.4 70.0 61.2 66.0 47.5
  • and denote the best result and second result, respectively.
Baselines. We combine Lamole into three BERT family models, such as DistilBert [48], DeBerta [49], and Bert to evaluate the performance of the proposed Lamole. For evaluating classification accuracy, we compare our Lamole with one SMILES string-based LM, ChemBERTa [10] and several GNNs including GCN [41], DGCNN [42], edGNN [43], GIN [44], RW-GNN [45], DropGNN [46], and IEGN [47].
For evaluating explanation accuracy, Lamole is compared with three types of alternative methods: 1) GCN with feature-based explainability techniques, including SmoothGrad [50], CAM [51], GradInput [52], and GradCAM [13], 2) Bert with the above feature-based explainability techniques, where Group SELFIES is used as input for a fair comparison, and 3) explainable GNNs including OrphicX [11], GNNExplainer [12], PGExplainer [21], and Gem [20].
Implementation Details. Following ChemBERTa [10], our used BERT family models are pre-trained on 100,000 molecules from the ZINC dataset [53]. 15% tokens in each input string are randomly for masked language learning. We finetuned the models on Mutag, Mutagen, PTC-FM, PTC-FR, PTC-MM, PTC-MR, and Liver. For each dataset, the ratio of samples with annotation masks over the data size is 10%. The details of experimental settings can be found in Appendix 3.
4.2 Results
Table 2: Mean Explanation Accuracy on the Seven Datasets (%)
Methods Mutag Mutagen PTC-FM PTC-FR PTC-MM PTC-MR Liver
SmoothGrad+GCN [50] 68.0 66.1 66.8 62.2 64.5 61.9 66.0
CAM+GCN [51] 66.8 65.4 66.9 62.7 61.4 65.5 67.5
GradInput+GCN [52] 69.1 65.1 68.5 64.3 62.3 62.0 67.2
GradCAM+GCN [13] 68.3 65.7 70.4 65.4 65.8 66.6 68.4
SmoothGrad+Bert [50] 71.3 71.3 72.8 70.6 65.2 68.9 74.4
CAM+Bert [51] 70.1 69.8 66.3 69.4 63.9 66.2 72.8
GradInput+Bert [52] 73.1 71.1 72.1 66.6 64.0 68.1 75.0
GradCAM+Bert [13] 73.5 71.5 76.3 69.0 69.1 71.2 75.8
OrphicX [11] 76.4 70.4 71.1 68.4 69.0 69.1 73.0
GNNExplainer [12] 69.5 63.0 68.6 67.3 66.0 66.3 71.4
PGExplainer [21] 65.7 58.0 69.3 66.3 64.7 65.1 71.0
Gem [20] 72.8 65.5 69.8 66.8 68.2 68.3 72,5
Lamole+DistilBert 70.9 73.0 74.0 70.2 69.6 78.1 76.1
Lamole+DeBerta 76.1 75.0 79.9 72.1 70.3 77.2 75.0
Lamole+Bert 77.8 75.2 81.1 72.2 72.0 73.1 77.3
Figure 4: Explanation visualization of one molecule (ID: 155) from the Mutag dataset.
(a) Mutag
(b) PTC-FM
(c) PTC-MR
Figure 5: The explanation accuracy of Lamole with different annotation rates on the Mutag, PTC-FM, and PTC-MR datasets.
(a) Mutag
(b) PTC-FM
(c) PTC-MR
Figure 6: The explanation plausibility of the compared algorithms on the Mutag, PTC-FM, and PTC-MR datasets.
Molecular Property Prediction. Table 1 shows the classification accuracy of compared algorithms. Our proposed Lamole+DistilBert, Lamole+DeBerta, and Lamole+Bert slightly outperform most compared algorithms on the Mutag, PTC-FM, PTC-FR, PTC-MR, and Liver datasets. Though the LMs, including Lamole and ChemBERTa, demonstrate good performance for molecular property prediction, they are inferior to GNN-based methods on the Mutagen and PTC-MM datasets. In addition, Lamole models show superior performance over ChemBERTa. This suggests that using molecular representations with more chemical semantics, like Group SELFIES, can help LMs better learn the structure-property relationships.
Explanation Performance. Table 2 presents the explanation accuracy of the compared explainability techniques. It is obvious that the LM-based methods using the Group SELFIES strings demonstrate significantly higher explanation accuracy over explainable GNNs. Specifically, Lamole improves the explanation accuracy by $1.4\%\sim 14.8\%$ compared to the baseline methods. This is because the molecular graph-based explainable GNNs can only highlight individual atoms and bonds rather than chemically meaningful substructures like functional groups. Instead, Group SELFIES explicitly encode the chemically meaningful substructures as tokens, and using Group SELFIES can help explainable methods to generate chemically meaningful explanations by highlighting the corresponding tokens.
We also selected some representative molecules for explanation visualization. These explanations are shown in Figs. 1, 4, and 12$\sim$15 (see Appendix), respectively. The right panel of those figures is the importance scores obtained by Lamole, where "other" in the figures is the average importance score of other unlisted functional groups/fragments. As shown in Fig. 4, our proposed Lamole assigns the highest two importance scores to the two ground truth function groups, i.e., benzene-1 and nitro group. In contrast, OrphicX cannot provide chemically meaningful explanations, and CradCAM+Bert cannot accurately identify the functional group interactions. Compared to baseline methods, Lamole provides chemically meaningful explanations. Particularly, the interaction among the functional groups is successfully captured. The visualization demonstrates Lamole’s superior interpretation in uncovering the structure-property relationships.
We also investigated the explanation accuracy of our Lamole under different ground truth annotation rates (10%, 20%, 50%, and 100%), as presented in Figs. 5 and 10 (see Appendix). Clearly, using more molecules with ground truth annotations can continuously improve explanation accuracy. In addition, compared to the results without using the marginal loss $\mathcal{L}_{M}$, using only 10% molecules with ground truth annotations can significantly improve explanation accuracy by up to 5%, confirming the effectiveness of using annotations. Though increasing the annotated ground truth rate can only bring a limited increase in the explanation accuracy, there is a trade-off between the achieved explanation accuracy and the additional annotation costs.
In addition, we evaluated the performance of compared algorithms by using the proposed explanation plausibility metric $\mathrm{EP}$. The statistical results of $\mathrm{EP}$ are presented in Figs. 6 and 11 (see Appendix). From the figures, we can observe that the $\mathrm{EP}$ values of the comparison algorithm are slightly lower, which means that the algorithms cannot confidently reflect the relationships between structure and property. Compared to the comparison algorithm, the $\mathrm{EP}$ values of Lamole have increased by $2\%\sim 9\%$. The statistical results of $\mathrm{EP}$ in Figs. 6 and 11 (see Appendix) indicate Lamole can provide explanations with higher confidence in matching the ground truth substructures.
Due to the page limitation, more analysis regarding the ablation studies on attention weights and the marginal loss, the impact of annotation rates, the analysis regarding the explanation plausibility, and more visualization results can be found in the Appendix.
5 Conclusions
This work developed a new framework for explainable molecular property prediction based on language models, called Lamole. Lamole uses Group SELFIES as input for perceiving the chemically meaningful semantics. By disentangling the information flows of Transformer-based LMs, Lamole integrates attention weights into gradients to generate explanations that can capture the functional group interactions. Furthermore, one marginal loss is designed to calibrate the explanations by aligning them with the chemists’ annotation. With the above designs, Lamole can provide chemically meaningful explanations while faithfully uncovering the structure-property relationships. Lamole ’s effectiveness has been demonstrated through theoretical analysis and extensive experimental validation.